Use the slope intercept form of the linear equation to write the equation of each line with the given slope and y-intercept.
Slope -1; y-intercept (0,1)
Please help!

Answer by Alan3354(30993)

Use the slope intercept form of the linear equation to write the equation of each line with the given slope and y-intercept.
Slope -1; y-intercept (0,1)
-------------
Use y = mx+b
y = -x + 1
Or, x+y = 1